OpenStudy (anonymous):

I'm so confused T.T help needed please~ To skate at Roller Heaven, each person must pay a membership fee (m) and a fee for each session (f). Art attended 5 sessions and paid $25 in all. Later, when he had attended 10 sessions, he paid $40 in all. How much is the membership fee and how much are the sessions. Show how to solve this, using any method.

OpenStudy (studyhard):

I solved it! Typing out solution now.

OpenStudy (anonymous):

oh thank you!! ^^ ~

OpenStudy (studyhard):

Solve this question using simultaneous equations. Let membership fee be x. Let fee for each session by y. First, he attended 5 sessions (1 membership fee + 5 session fees). x+5y=25 (first equation). Next, he attended another 10 sessions (1 membership fee + 10 session fees). x+10y=40 (second equation). If we take the second equation and minus the first equation, x-x+10y-5y=40-25. 5y=15. y=3. Therefore, each session fee would be $3. Next, we substitute this value of y into equation 1. x+5y= 25. x+3(5)=25. x+15=25. x=10. OVERALL SUMMARY: Membership fee: $10. Session fee: $3 each

OpenStudy (studyhard):

No prob! :) hope i did the question correctly :/

OpenStudy (anonymous):

thank you! i was so stumped ^^ !

OpenStudy (studyhard):

Hahaha don't mention it! I just hope you understood what I'm saying.